# Data-Efficient RLVR via Off-Policy Influence Guidance

### v1

@@ E.3 Baselines in Main Results
For the above baselines, data selection is performed once globally using the initial policy $\pi_{\theta_{0}}$ at a selection ratio of $\alpha=0.1$. We also include DAPO (Yu et al., 2025), a batch-level filtering method that removes samples with perfect or zero pass rates and replaces them from a buffer of historical trajectories.
E.4 Prompt Templates for Evaluation & RL Training
@@ E.5 Hyperparameters of RL
Appendix F Additional Analysis on Off-policy Gradient Estimator
@@
Appendix G Additional Analysis on Gradient Projection
@@
Appendix H Additional Analysis on Data Selected by CROPI
@@
H.1 Data Source of Selected Prompts
@@
H.2 Knowledge Domains of Selected Prompts
@@
Since direct inspection of each individual problem makes it difficult to observe distinctive patterns, we adopted the categorization approach from s1 Muennighoff et al. (2025). Utilizing GPT-4o, we classified both the validation set (randomly sampled 100 questions) and the top-100 mathematics problems selected by POPI according to the Mathematics Subject Classification (MSC) system (e.g., geometry, combinatorics, etc.) from the American Mathematical Society [^2] https://mathscinet.ams.org/mathscinet/msc/msc2020.html. The results (Figure 12 show that the distribution of knowledge domains in the selected data closely mirrors that of the validation set itself. Moreover, as training progresses, the distribution of knowledge domains among the selected samples also shifts dynamically to meet the need of online policy.
@@
H.3 Semantic Diversity of Selected Prompts
@@
H.4 Human Annotated Difficulty-level of Selected Prompts
@@
H.5 Most influential questions
@@
Appendix I Additional Results on CROPI Experiments

### v2

@@ E.3 Baselines in Main Results
For the above baselines, data selection is performed once globally using the initial policy $\pi_{\theta_{0}}$ at a selection ratio of $\alpha=0.1$. We also include DAPO (Yu et al., 2025), a batch-level filtering method that removes samples with perfect or zero pass rates and replaces them from a buffer of historical trajectories. All reported Acc.@step metrics in the main text are computed from evaluation curves smoothed by a moving average with window size $5$.
E.4 Additional Convergence and Time-based Comparison
To better characterize the final-stage behavior of CROPI, we extended the 1.5B experiment beyond the initial 1k-step budget and compared the best smoothed targeted accuracy achieved within 2000 training steps. The results in Table 3 show that CROPI reaches its peak performance earlier and attains a slightly higher best accuracy than standard GRPO within the same step budget.
Table 3: Peak targeted accuracy within 2000 training steps on the 1.5B setting. Accuracy is computed from curves smoothed with a moving average of window size $5$.
Method Step Peak Accuracy (%)
CROPI (Ours) 1130 71.05
Standard GRPO 1790 70.17
We also evaluated both methods under fixed wall-clock budgets. As shown in Table 4, CROPI consistently outperforms standard GRPO at 12h, 18h, and 24h, which complements the step-level speedup reported in the main text.
Table 4: Targeted accuracy (%) under fixed wall-clock budgets for the 1.5B setting.
Method 12h 18h 24h
Standard GRPO 67.73 68.84 69.34
CROPI (Ours, 10%) 68.65 69.73 69.91
E.5 Random Selection Baseline
To isolate the contribution of influence-based selection itself, we additionally evaluate a phase-wise random baseline that keeps the CROPI training schedule unchanged but replaces influence ranking with random selection of 10% of the training data at each phase. Table 5 shows that random phase-wise selection underperforms both full-data GRPO and CROPI, indicating that the gain does not come from simply reducing the training set size.
Table 5: Ablation with phase-wise random selection on the 1.5B setting.
Method Acc.@500 Acc.@1000
Full Data + GRPO 68.04 69.42
CROPI w. Random Selection (10%) 67.29 68.04
CROPI (Ours) 69.49 70.26
E.6 Prompt Templates for Evaluation & RL Training
@@ E.7 Hyperparameters of RL
Appendix F Additional Analysis on Rollout Cost
To clarify why rollout-free influence estimation is important in RLVR, we compare the cost of a rollout step against a single forward+backward step in the 1.5B setting. For a fair comparison, both settings use batch size 128 with sequence length capped at 8192. Table 8 shows that rollout generation is substantially more expensive than gradient computation under comparable settings.
Table 8: Cost comparison between rollout and forward+backward in the 1.5B setting.
Operation Avg. Time (s) Relative Cost
Forward+Backward 35.91 1.00x
Rollout 319.80 8.91x
This difference is also aligned with the computational characteristics of the two operations. A forward+backward step mainly consists of dense tensor computation and is typically compute-bound, while autoregressive rollout repeatedly reads model weights during decoding and is often constrained by memory bandwidth. As a result, methods that require fresh rollouts for every candidate prompt incur a much larger selection overhead than methods that reuse offline trajectories.
Appendix G Additional Analysis on Off-policy Gradient Estimator
@@
Appendix H Additional Analysis on Gradient Projection
@@
Appendix I Additional Analysis on Data Selected by CROPI
@@
I.1 Data Source of Selected Prompts
@@
I.2 Knowledge Domains of Selected Prompts
@@
Since direct inspection of each individual problem makes it difficult to observe distinctive patterns, we adopted the categorization approach from s1 Muennighoff et al. (2025). Utilizing GPT-4o, we classified both the validation set (randomly sampled 100 questions) and the top-100 mathematics problems selected by POPI according to the Mathematics Subject Classification (MSC) system (e.g., geometry, combinatorics, etc.) from the American Mathematical Society [^2] https://mathscinet.ams.org/mathscinet/msc/msc2020.html. The results (Figure 12) show that the distribution of knowledge domains in the selected data closely mirrors that of the validation set itself. Moreover, as training progresses, the distribution of knowledge domains among the selected samples also shifts dynamically to meet the need of online policy.
@@
I.3 Semantic Diversity of Selected Prompts
@@
I.4 Human Annotated Difficulty-level of Selected Prompts
@@
I.5 Most influential questions
@@
Appendix J Additional Results on CROPI Experiments